OpenStudy (lilshane):

What is the ratio 5 to 105 expressed as a fraction in simplest form?

OpenStudy (sierra!):

What do you think again?

OpenStudy (sierra!):

Were pretty much doing the same thing just with different numbers

OpenStudy (lilshane):

its weird idk 5/5

OpenStudy (lilshane):

@iambatman

OpenStudy (lilshane):

@pooja195

OpenStudy (lilshane):

@NeonStrawsForever

OpenStudy (lilshane):

would it be 5/5

OpenStudy (lilshane):

@NeonStrawsForever

OpenStudy (sierra!):

I think its \[\frac{ 1 }{ 12 }\]

OpenStudy (sierra!):

sorry

OpenStudy (sierra!):

\[\frac{ 1 }{ 21 }\]

OpenStudy (lilshane):

@iGreen

OpenStudy (lilshane):

\[\frac{ 5 }{ 5 }\]

OpenStudy (jordan123321):

i say 5/10

OpenStudy (lilshane):

actually i think 1/21

OpenStudy (anonymous):

You're right.

OpenStudy (lilshane):

thanks guys i got 100%

OpenStudy (anonymous):

You're welcome ^_^

OpenStudy (lilshane):

wanna give me a medal lol

OpenStudy (lilshane):

@NeonStrawsForever

OpenStudy (anonymous):

Yes. Since you got it right on your first try.

OpenStudy (lilshane):

scince i anwser it even though i had some help